OpenStudy (anonymous):

A bag contains 7 lavender marbles, 11 turquoise marbles and 5 rainbow marbles. What is the chance of drawing a turquoise marble? If a turquoise marble is drawn then placed back into the bag, and a second marble is drawn, what is the probability of drawing a rainbow marble? Give solutions exactly in reduced fraction form, separated by a comma.

OpenStudy (anonymous):

There is an 11/23 chance of picking a turquoise marble from the whole lot the first time. There is a 5/23 chance that the second marble drawn is rainbow.\[{11\over23},{5\over23}\]

OpenStudy (anonymous):

total=7+11+5=23 p(turquoise marble) = turquoise marbles/total marbles =11/23 p(rainbow marble) = rainbow marbles/total marbles =5/23 11/23,5/23

OpenStudy (anonymous):

as a matter of interest the chance of getting a turquoise then a rainbow is obtained by multiplying the 2 results i.e. 11/23 * 5/23